Suppose a monopolist faces a market demand of P = 48 - 4Q. The monopolist has a constant marginal cost of 8 per unit. If the monopolist can only charge a single price to consumers, how many units should the monopolist produce to maximize profits?

10

5

12

6
Answer #1

Monopoly

P=48-4Q

TR=PQ

=(48-4Q)Q

=48Q-4Q2

MR=dTR/dQ

=48-8Q

MC=8

Monopoly profit maximizing condition is

MR=MC

8=48-8Q

8Q=48-8

Q=40/8

Q=5

Hence option second is the correct answer.